# Entity-based Reinforcement Learning for Autonomous Cyber Defence

### v1

@@ Abstract
A significant challenge for autonomous cyber defence is ensuring a defensive agent’s ability to generalise across diverse network topologies and configurations. This capability is necessary for agents to remain effective when deployed in dynamically changing environments, such as an enterprise network where devices may frequently join and leave. Standard approaches to deep reinforcement learning, where policies are parameterised using a fixed-input multi-layer perceptron (MLP) expect fixed-size observation and action spaces. In autonomous cyber defence, this makes it hard to develop agents that generalise to environments with network topologies different from those trained on, as the number of nodes affects the natural size of the observation and action spaces. To overcome this limitation, we reframe the problem of autonomous network defence using entity-based reinforcement learning, where the observation and action space of an agent are decomposed into a collection of discrete entities. This framework enables the use of policy parameterisations specialised in compositional generalisation. Namely, we train a Transformer-based policy on the Yawning Titan cyber-security simulation environment and test its generalisation capabilities across various network topologies. We demonstrate that this approach significantly outperforms an MLP-based policy on fixed networks, and has the ability for zero-shot generalisation to networks of a different size to those seen in training. These findings highlight the potential for entity-based reinforcement learning to advance the field of autonomous cyber defence by providing more generalisable policies capable of handling variations in real-world network environments.
@@ 2.2 Entities and Objects in Reinforcement Learning
In this section we provide an overview of attempts to leverage the underlying structure of RL environments, as well as practical approaches for dealing with complex environments with potentially variable observation and action spaces, such as in the case of autonomous cyber defence. In the first instance, particularly in the case of Factored MDPs Boutilier et al., (2000), the motivation is to find more efficient solution methods by exploiting the ‘factorised’ structure of an environment. More broadly, all of these approaches contribute to solving the problem of compositional generalisation (Zhao et al.,, 2022; Lin et al., 2023b, ). That is, training policies that are able to generalise across environments containing different compositions of familiar objects or entities and dynamics. For further reference, Mohan et al., (2024) provide a comprehensive survey of structure in reinforcement learning.
@@ 2.2.3 Relational Reinforcement Learning
In RRL, the state space, action space, and transition function are represented using relational descriptions. This allows for a more compact and expressive representation of complex environments. The key idea is to learn relational policies that can generalise across objects and relations, rather than learning separate policies for each specific instance van Otterlo, (2005).
RRL has been applied to various domains, including robotics walker2004relational, game playing guestrin2003generalizing, and natural language processing branavan2012learning. Its ability to handle structured environments and transfer knowledge across tasks makes it a promising approach for developing more generalisable and sample-efficient reinforcement learning algorithms.
@@ 2.2.5 Graph-based Reinforcement Learning for robotic control
2.2.6 Generalisability in Single-Agent RL
Finally, we discuss some relevant popular work on context generalisability in single-agent RL Kirk et al., (2023). Within this domain, contributions such as Zhang et al., (2020); Seitzer et al., (2021) and Han et al., (2021); Ghosh et al., (2021) have focused on developing methods that expose the agent to a broad range of contextual episodic variability during training, enabling it to learn robust and invariant causal representations across different ‘contexts’ (mostly across episodes). These policies are then proven to generalise better to new, unseen contexts at test time. The difference between these methods and the entity-based approach used in this work is mainly conceptual and lies in the distinction between the time-invariant context and the entity-level representations that can vary within-episode, as they are treated as part of the evolving environment’s state space.
@@ 3.1 Entity Gym
For the GlobalCategoricalActionSpace, one might assign actions that affect a large number of entities simultaneously, or have no direct effect on the entities in the environment but change the state of the agent. For example, moving the top-down camera in a video game such as StarCraft Vinyals et al., (2019), which changes the entities that are visible to the agent. For an autonomous network-based agent, this could be any action that has network-wide implications, or perhaps the decision to intervene in the network if malicious activity is detected Hammar and Stadler, (2022).
@@ 6.1 Training
All of our experiments employed the Yawning Titan configuration described in Section 4, with variable randomised network topologies of 10, 20 and 40 nodes generated via the Erdős-Renyi method. We trained each agent on 10,000,000 total environment steps, which amounts to 10,000 episodes with a fixed length of 100 timesteps per episode. During training, we evaluated each policy every 10,000 environment steps and logged the resulting episodic reward to produce training graphs. Exchangeability. As suggested in Mern et al., (2020, 2019), one advantage of the entity-based approach and using a Transformer policy is ‘object exchangeability’. That is, the robustness of Transformers and GNNs to different permutations of objects in the observation vector or action space without the effectiveness of the policy being impacted. To demonstrate this property, we consider two different training settings, which we refer to as Static and Random. In the Static setting, a single network is randomly initialised at the start of training, and remains the same for every episode; the only thing that varies is the identity of the entry node. In the Random setting, an entirely new network is initialised for every episode. The aim of this setup is to evaluate policy robustness across inhomogeneous networks in the training phase, and is similar to the method used by Hong et al., (2022) in ‘Multitask Inhomogeneous Reinforcement Learning’ for robotic control.
@@
Figure 2(a) shows the reward trajectories of all four types of policies trained on 10-node networks. Both entity-based agents (‘Entity_random_10’ and ‘Entity_static_10’) show similar reward trajectories that converge rapidly and plateau higher than their counterpart Stable Baselines MLP-based policies, close to the maximum possible reward of 100. The Stable Baselines PPO agents trained on the single static network (‘sb3_static_10’) converged quickly as well, but to a lower level of asymptotic rewards than the entity-based agents. In line with expectations, the baseline PPO agents trained on a series of random networks that vary across episodes (‘sb3_random_10’) struggled to learn a good policy and displayed high variance in their performance. This supports the notion that a Transformer entity-based policy is more robust to exchangeable representations, and performs no worse than when trained on a static environment, whereas a fixed MLP-based policy struggles to learn an effective policy when the object that each element of the agent’s observation vector represents changes with each episode (Mern et al.,, 2020, 2019; Zhao et al.,, 2022).
Figures 2(b) and (c) show similar results on 20-node and 40-node networks respectively. Here, the difference between the episodic reward of the two entity-based policies and the statically trained Stable Baselines PPO agent is less significant. This may be, in part, due to the nature of the reward function defined as the proportion of non-compromised nodes at time $t$. Since the skill and behaviour of the red agent are unchanged, this naturally means that a lower proportion of nodes will be compromised on a given timestep, and thus rewards for the blue agent will be higher in larger environments. Nevertheless, the Stable Baselines agent trained on randomly generated networks still does comparatively poorly, and struggles to learn a policy that is invariant to different network topologies.
@@ 7 Discussion
In this paper, we have demonstrated the basic functionality and the strong generalisation capabilities of an entity-based RL method through the use of the Entity Gym framework for the Yawning Titan autonomous cyber defence environment. We suggest that other, similar or more complex cyber defence simulators could be readily adapted in the same way, due to the fundamental node-based structure of the observation and action spaces. In fact, when designing a simulator, it is likely to be much more straightforward to use an interface such as Entity Gym than devising rules for how a fixed-input agent can interact through the standard Gymnasium interface. One drawback is the lack of out-of-the-box implementations of algorithms designed to work with entity-based environments, and tuned for optimal performance with complex Transformer or GNN policies. Indeed, as can be seen from the training plots, the RogueNet policy displays occasional instability during training in the environment setting we used. However we believe this is outweighed by the potential for training and generalisation across variable network topologies, and can be rectified with more development and testing. We suggest that future cyber defence environments be designed using an ‘entity-based’ philosophy, and constructed using an interface akin to Entity Gym.

### v2

@@ Abstract
A significant challenge for autonomous cyber defence is ensuring a defensive agent’s ability to generalise across diverse network topologies and configurations. This capability is necessary for agents to remain effective when deployed in dynamically changing environments, such as an enterprise network where devices may frequently join and leave. Standard approaches to deep reinforcement learning, where policies are parameterised using a fixed-input multi-layer perceptron (MLP) expect fixed-size observation and action spaces. In autonomous cyber defence, this makes it hard to develop agents that generalise to environments with network topologies different from those trained on, as the number of nodes affects the natural size of the observation and action spaces. To overcome this limitation, we reframe the problem of autonomous network defence using entity-based reinforcement learning, where the observation and action space of an agent are decomposed into a collection of discrete entities. This framework enables the use of policy parameterisations specialised in compositional generalisation. We train a Transformer-based policy on the Yawning Titan cyber-security simulation environment and test its generalisation capabilities across various network topologies. We demonstrate that this approach significantly outperforms an MLP-based policy when training across fixed-size networks of varying topologies, and matches performance when training on a single network. We also demonstrate the potential for zero-shot generalisation to networks of a different size to those seen in training. These findings highlight the potential for entity-based reinforcement learning to advance the field of autonomous cyber defence by providing more generalisable policies capable of handling variations in real-world network environments.
@@ 2.2 Entities and Objects in Reinforcement Learning
In this section we provide an overview of attempts to leverage the underlying structure of RL environments, as well as practical approaches for dealing with complex environments with potentially variable observation and action spaces, such as in the case of autonomous cyber defence. In the first instance, particularly in the case of Factored MDPs (Boutilier et al.,, 2000), the motivation is to find more efficient solution methods by exploiting the ‘factorised’ structure of an environment. More broadly, all of these approaches contribute to solving the problem of compositional generalisation (Zhao et al.,, 2022; Lin et al., 2023b, ). That is, training policies that are able to generalise across environments containing different compositions of familiar objects or entities and dynamics. For further reference, Mohan et al., (2024) provide a comprehensive survey of structure in reinforcement learning.
@@ 2.2.3 Relational Reinforcement Learning
In RRL, the state space, action space, and transition function are represented using relational descriptions. This allows for a more compact and expressive representation of complex environments. The key idea is to learn relational policies that can generalise across objects and relations, rather than learning separate policies for each specific instance (van Otterlo,, 2005).
@@ 3.1 Entity Gym
For the GlobalCategoricalActionSpace, one might assign actions that affect a large number of entities simultaneously, or have no direct effect on the entities in the environment but change the state of the agent. For example, moving the top-down camera in a video game such as StarCraft (Vinyals et al.,, 2019), which changes the entities that are visible to the agent. For an autonomous network-based agent, this could be any action that has network-wide implications, or perhaps the decision to intervene in the network if malicious activity is detected (Hammar and Stadler,, 2022).
@@ 6.1 Training
All of our experiments employed the Yawning Titan configuration described in Section 4, with variable randomised network topologies of 10, 20 and 40 nodes generated via the Erdős-Renyi method. We trained each agent on 10,000,000 total environment steps, which amounts to 10,000 episodes with a fixed length of 100 timesteps per episode. During training, we evaluated each policy every 10,000 environment steps and logged the resulting episodic reward to produce training graphs.
Exchangeability.
As suggested in Mern et al., (2020, 2019), one advantage of the entity-based approach and using a Transformer policy is ‘object exchangeability’. That is, the robustness of Transformers and GNNs to different permutations of objects in the observation vector or action space without the effectiveness of the policy being impacted. To demonstrate this property, we consider two different training settings, which we refer to as Static and Random. In the Static setting, a single network is randomly initialised at the start of training, and remains the same for every episode; the only thing that varies is the identity of the entry node. In the Random setting, an entirely new network is initialised for every episode. The aim of this setup is to evaluate policy robustness across inhomogeneous networks in the training phase, and is similar to the method used by Hong et al., (2022) in ‘Multitask Inhomogeneous Reinforcement Learning’ for robotic control.
Hyperparameters
To produce the training graphs we used the default PPO hyperparameters as specified in the libraries for both agents. It’s worth noting that the default learning rate for the optimiser for the entity-based policy is 0.005, compared to 0.0003 in the Stable Baselines 3 implementation. Different choices of learning rate are likely to affect the gap between asymptotic reward of the two varieties of policy in the Static training regime.
@@
Figure 2(a) shows the reward trajectories of all four types of policies trained on 10-node networks. Both entity-based agents (‘Entity_random_10’ and ‘Entity_static_10’) show similar reward trajectories that converge rapidly and plateau close to the maximum possible reward of 100. The Stable Baselines PPO agents trained on the single static network (‘sb3_static_10’) converged quickly as well, to a slightly lower asymptotic reward. In line with expectations, the baseline PPO agents trained on a series of random networks that vary across episodes (‘sb3_random_10’) struggled to learn a good policy and displayed high variance in their performance. This supports the notion that a Transformer entity-based policy is more robust to exchangeable representations, and performs no worse than when trained on a static environment, whereas a fixed MLP-based policy struggles to learn an effective policy when the object that each element of the agent’s observation vector represents changes with each episode (Mern et al.,, 2020, 2019; Zhao et al.,, 2022).
Figures 2(b) and (c) show similar results on 20-node and 40-node networks respectively. Here, the difference between the asymptotic episodic reward of the policies trained in the Static regime grows smaller with a larger number of nodes. This may be, in part, due to the nature of the reward function defined as the proportion of non-compromised nodes at time $t$. Since the skill and behaviour of the red agent are unchanged, this naturally means that a lower proportion of nodes will be compromised on a given timestep, and thus rewards for the blue agent will be higher in larger environments. Nevertheless, the Stable Baselines agent trained on randomly generated networks still does comparatively poorly, and struggles to learn a policy that is invariant to different network topologies.
@@ 7 Discussion
In this paper, we have demonstrated the basic functionality and the strong generalisation capabilities of an entity-based RL method through the use of the Entity Gym framework for the Yawning Titan autonomous cyber defence environment. We suggest that other, similar or more complex cyber defence simulators could be readily adapted in the same way, due to the fundamental node-based structure of the observation and action spaces. In fact, when designing a simulator, it is likely to be much more straightforward to use an interface such as Entity Gym than devising rules for how a fixed-input agent can interact through the standard Gymnasium interface. One drawback is the lack of out-of-the-box implementations of algorithms designed to work with entity-based environments, and tuned for optimal performance with complex Transformer or GNN policies. Indeed, as can be seen from the training plots, the RogueNet policy displayed occasional instability during training in the environment setting and hyperparameters we used, although this might rectifiable by using a lower learning rate. We believe any shortcomings in out-of-the-box performance of an ‘entity-based’ policy is outweighed by the potential for training and generalisation across variable network topologies, and can be rectified with more development and testing. We suggest that future cyber defence environments be designed using an ‘entity-based’ philosophy, and constructed with support for using an interface akin to Entity Gym.